# Random matrix universality in dynamical correlation functions at late times

### v2

@@ 1 Introduction
$\overline{\mathcal{C}}(t)=\mathcal{C}_{\text{d}}(t)+\mathcal{C}_{\mathrm{c}}(t)\ ,$ (2)
for a suitably-defined connected correlation function $\mathcal{C}(t)$ which depends on the RMT universality class of the system. We show that $\mathcal{C}_{\text{d}}(t)$ amounts for the non-universal physical dynamical correlations (i.e. the Fourier transform of the smooth ETH off-diagonal function), while $\mathcal{C}_{\mathrm{c}}(t)$ can be written as the convolution of the latter with the spectral correlations, and thus represents the spectral correlations responsible for the late-time ramp-plateau behaviour.
Similar observations have been drawn in the high-energy literature [7, 21, 22, 23, 24], random circuits [25] or hydrodynamics [26]. While the mechanism by which such a ramp can appear is easy to grasp, a justification of when it can be actually observed is instead rather subtle. Here, we provide a careful analytical study using the ETH approach. Our main assumption is that the average over energy levels and eigenvectors decouples. This is a standard result for rotationally-invariant Hamiltonians, and we here assume it also holds true for realistic many-body systems satisfying the ETH. We also show numerical evidence for our predictions in macroscopic many-body systems. While we chose to study numerically spin glass models, the same conclusions could have been drawn considering the SYK model, which however was not convenient for us as we need to treat separately different RMT universality classes. A key requirement to expect such a RMT behaviour is to consider observables whose expectation values (one-point functions) do not show any energy dependence. This allows, in absence of degenerate energy levels, the plateau to arise mainly from the fluctuations of the observable’s diagonal matrix elements, making it exponentially small in the system size. This is to be put in contrast with the case considered in [27, 28], and recently in [29], where an explicit energy dependence leads to a plateau which scales polynomially with the system size.
@@ 2.1 Two-time correlations and SFF in RMT
with the spectral form factor (SFF) defined as ${\rm SFF}(t)=\frac{1}{\mathcal{N}^{2}}\sum_{i,j}e^{i(E_{i}-E_{j})t}$. It is natural to introduce the shifted connected correlation function $\mathcal{C}$ such that
$\mathcal{C}(t)=C(t)-\left(1-\frac{\beta_{\rm RMT}}{2}\right)\lim_{t\to\infty}\overline{C}(t)\ ,$ (17)
@@
$\overline{\mathcal{C}}(t)=\kappa_{2}\overline{{\rm SFF}}(t)\ .$ (18)
This identity shows that, for a Hamiltonian drawn from random ensembles, the behaviour of a two-point function can be extracted from that of the ${\rm SFF}$ [7, 31]. The present analysis emphasises the importance of the symmetry classes for extracting the eigenvalue correlations from dynamical correlators, which is encoded in the re-scaled Eq.(17) and not in the bare connected correlator. From Eq.(18), it follows that, for random matrices, the decomposition of dynamical correlations in Eq.(2) holds upon identifying $\mathcal{C}_{\text{d}}(t)=\kappa_{2}{\rm SFF}_{\text{d}}(t)$ and $\mathcal{C}_{\mathrm{c}}(t)=\kappa_{2}{\rm SFF}_{\mathrm{c}}(t)$.
@@ 2.2 Lack of self-averaging
While the previous calculation characterised the ensemble-averaged quantity $\overline{\mathcal{C}}(t)$, this section aims at estimating the single instance fluctuations $\delta\mathcal{C}(t)=\mathcal{C}(t)-\overline{\mathcal{C}}(t)$. In random matrices, because $\overline{|A_{ij}|^{4}}\propto\overline{|A_{ij}|^{2}}^{2}$ with a proportionality constant depending on the ensemble, we expect that matrix elements behave as
@@
with $\{\xi_{ij}\}_{ij}$ a set of weakly correlated random variables with mean zero, variance one, and whose exact distributions depend on the observable considered. In a single instance $\mathcal{C}(t)$, summing the $\mathcal{N}^{2}$ noise contributions gives rise to a noisy component of strength $\mathcal{N}^{-1}$, i.e.
$\mathcal{C}(t)=\overline{\mathcal{C}}(t)+\mathcal{O}(\mathcal{N}^{-1})\ .$ (20)
While this does not significantly obstruct the recovery of $\overline{\mathcal{C}}(t)$ in the slope of order $\mathcal{O}(1)$ and the plateau of order $\mathcal{O}(\mathcal{N}^{-1})$, fluctuations completely hide the signal close to the dip which scales as $\mathcal{N}^{-3/2}$ (see Sec. 2.1).
This is confirmed by the numerics in the Gaussian ensemble case (see the right panels of Figs. 1-2 below) where we plot the single-sample $\mathcal{C}(t)$. Contrary to the Spectral Form Factor, where the small-window time average reproduces the dip-ramp behaviour, for the dynamical correlation function $\mathcal{C}(t)$, the dip and ramp are nowhere to be seen even after averaging over a small time window. The fact that, in the RMT toy models, the noise $\delta\mathcal{C}(t)$ can become larger than the average signal $\overline{\mathcal{C}}(t)$ thus results in correlation functions not being self-averaging, and in the small-window time-average not reproducing the ensemble-average in contrast to what is usually expected in ETH in the decaying part.
Figure 1: Left: $(a)$ Shifted correlation function $\overline{\mathcal{C}}(t)$ and $(b)$ spectral form factor $\overline{{\rm SFF}}(t)$ averaged over 1000 GOE Hamiltonians. Right: $(c)$ Shifted correlation function $\mathcal{C}(t)$ and $(d)$ spectral form factor ${\rm SFF}(t)$ obtained from a single GOE Hamiltonian. Plots $(c)$ and $(d)$ are for $L=12$ spins. In all plots, the thin curves are the exact result while the thick curves are a smoothed version obtained by convoluting with a Gaussian kernel. The horizontal dashed lines are the expected plateau values coming from the diagonal ensemble. The increasing oblique dashed lines are fits tproportional-$\propto t$.
@@ 2.3.1 The Gaussian Orthogonal Ensemble
We consider here a Hamiltonian from the GOE, i.e. drawn from Eq. (3) with $\beta_{\rm RMT}=1$, and take the observable $A=S^{z}_{j}$, with $j$ an arbitrary site index. Fig. 1 (right) shows the behaviour of $\mathcal{C}(t)=C(t)-\kappa_{2}/\mathcal{N}$ and of the SFF for a single realisation of the GOE Hamiltonian. Upon averaging over some small time window, one cannot resolve the ramp in $\mathcal{C}(t)$. However, if one considers the average $\overline{\mathcal{C}}(t)$ over many realisations of the GOE Hamiltonian, the slope-dip-ramp-plateau becomes clear as shown in Fig. 1 (left). This is in contrast with the SFF where, despite the fluctuations, the ramp is visible also in a single instance, and the time average of one single instance coincides with the ensemble average.
@@ 2.3.2 The Gaussian Unitary ensemble
We now consider a Hamiltonian from the GUE (Eq. (3) with $\beta_{\rm RMT}=2$). Fig. 2 (right) shows the behaviour of $\mathcal{C}=C(t)$ and of the SFF for a single realisation of the GUE Hamiltonian. Upon averaging over some small time window, the slope-dip-ramp-plateau feature becomes visible in the SFF, while the ramp cannot be resolved in $\mathcal{C}$. If one averages $\mathcal{C}$ and the SFF over many realisations of the GUE Hamiltonian, the ramp-plateau becomes much more clear as shown in Fig. 2 (left).
Figure 2: Left: $(a)$ Shifted correlation function $\overline{\mathcal{C}}(t)$ and $(b)$ spectral form factor $\overline{{\rm SFF}}(t)$ averaged over 1000 GUE Hamiltonians. Right: $(c)$ Shifted correlation function $\mathcal{C}(t)$ and $(d)$ spectral form factor ${\rm SFF}(t)$ obtained from a single GUE Hamiltonian. Plots $(c)$ and $(d)$ are for $L=12$ spins. In all plots, the thin curves are the exact result while the thick curves are a smoothed version obtained by convoluting with a Gaussian kernel. The horizontal dashed lines are the expected plateau values coming from the diagonal ensemble. The increasing oblique dashed lines are fits tproportional-$\propto t$.
@@ 2.3.3 Numerical evaluation of the fluctuations
In order to quantify the fluctuations of correlations relative to their average in Eq.(20), we consider the variance of $\mathcal{C}(t)$ normalised by its average, namely
$\overline{\delta\mathcal{C}^{2}}(t)/\overline{\mathcal{C}}^{2}(t),$ (21)
@@
where $\delta{\rm SFF}(t)={\rm SFF}(t)-\overline{{\rm SFF}}(t)$ are the fluctuations of the SFF. The results for the GOE case are summarised in Fig. 3, and a similar picture can be drawn for the GUE case. While the ${\rm SFF}$ has a bounded noise-to-signal ratio, for the correlation function $\mathcal{C}$, this ratio increases with the system size $L$ in the dip and ramp, signalling there that the fluctuations $\delta\mathcal{C}(t)$ completely dwarf the ensemble average $\overline{\mathcal{C}}(t)$.
@@
Figure 3: Normalised fluctuations of the shifted correlation function $\mathcal{C}(t)$ and the spectral form factor ${\rm SFF}(t)$ averaged over 1000 GOE Hamiltonians at inverse temperature $\beta=0$. The thin curves are the exact result while the thick curves are a smoothed version obtained by convoluting with a Gaussian kernel.
@@ 3 Many-body Hamiltonian
$\overline{\mathcal{C}}(t)=\mathcal{C}_{\text{d}}(t)+\mathcal{C}_{\mathrm{c}}(t)\ ,$ (23)
@@
$\mathcal{C}_{\text{d}}(t)=\int\text{d}\omega\,|f_{e_{0}}(\omega)|^{2}e^{i\omega t}\ ,$ (24)
where $f_{e_{0}}(\omega)$ is the smooth function appearing in the off-diagonal ETH ansatz (see below). On the other hand, the second contribution $\mathcal{C}_{\mathrm{c}}$ is related to the connected spectral correlations and generically features a late-time ramp followed by a plateau as for the SFF since
$\mathcal{C}_{\mathrm{c}}(t)\sim\int\text{d}t^{\prime}C_{\text{d}}(t-t^{\prime}){\rm SFF}_{\mathrm{c}}(t^{\prime})\simeq f^{2}_{e_{0}}(0){\rm SFF}_{\mathrm{c}}(t)\ .$ (25)
@@ 3.1 Two-time correlations and SFF within ETH
An important assumption made in the following is that the observable $A$ has an average ${\mathcal{A}}(e)$ that does not depend on $e$, while $A$ still fluctuates as in Eq. (29). This is, for instance, the case of a local observable protected by some global conservation law in a disordered system. This hypothesis is required to avoid extra terms which could hide the slope-dip-ramp-plateau behaviour we are looking for [29], because of polynomially small/large corrections to the plateau, which may be understood by the fact that the fluctuations of the observable are then driven by the large local excursions of energy.
Drawing inspiration from the previous section on RMT, we now wish to establish a link between the large-time behaviour of connected correlation functions defined in Eq. (13) and the ${\rm SFF}$. We assume for simplicity that the system is put at infinite temperature ($\beta=0$) since using a finite temperature is a straightforward generalisation. In the limit of large, yet finite, system size, the ensemble-averaged correlation function reads
@@
where $E=(E_{1}+E_{2})/2$ and $\omega=E_{1}-E_{2}$. Due to the standard saddle-point argument of the ETH, the contribution of the canonical average $\langle A\rangle^{2}$ exactly cancels that of the microcanonical average $\mathcal{A}(e)$. Shuffling terms between both integrals gives
@@
$\mathcal{C}(t)=C(t)-\left(1-\frac{\beta_{\rm RMT}}{2}\right)\lim_{t\to\infty}C(t)\ .$ (33)
@@
with $\frac{1}{L}\frac{\text{d}^{2}s}{\text{d}e^{2}}=-\beta^{2}/C_{V}$ and $C_{V}$ the heat capacity. This gives a contribution $\mathcal{C}_{\text{d}}$ to the shifted correlation function $\mathcal{C}$ defined by
$\mathcal{C}_{\text{d}}(t)=\frac{1}{Z}\int\text{d}E\,\text{d}\omega\,\overline{\rho(E)}e^{\frac{1}{4L}\frac{\text{d}^{2}s(e)}{\text{d}e^{2}}\omega^{2}}f^{2}_{e}(\omega)e^{i\omega t}\ .$ (35)
As in the RMT case, $\mathcal{C}_{\text{d}}$ is expected to give rise to the slope since it comes from the disconnected part of the correlations between energy levels. It is furthermore expected to be self-averaging with respect to the ensemble average, as is usual within the ETH. In the large system size limit, the factor $\exp\big{(}\frac{1}{4L}\frac{\text{d}^{2}s(e)}{\text{d}e^{2}}\omega^{2}\big{)}$ can be dropped and a saddle point approximation (also done for the partition function $Z$) leads to
${\mathcal{C}}_{\text{d}}(t)=\int\text{d}\omega\,f^{2}_{e_{0}}(\omega)e^{i\omega t}\ ,$ (36)
where $e_{0}$ is such that $\frac{\text{d}s}{\text{d}e}(e_{0})=0$.
@@
$\mathcal{C}_{\mathrm{c}}(t)=$ $\frac{1}{Z}\int\text{d}E\,\text{d}\omega\,\pi^{2}\frac{\overline{\rho(E+\omega/2)}\,\overline{\rho(E-\omega/2)}}{\overline{\rho(E)}}\mathcal{R}\big{[}\mathcal{N}(\Phi(E+\omega/2)-\Phi(E-\omega/2))\big{]}f^{2}_{e}(\omega)e^{i\omega t}$
@@
$\mathcal{C}_{\mathrm{c}}(t)\simeq\pi^{2}\int\text{d}\omega\,\mathcal{R}[\pi\overline{\rho(E_{0})}\omega]f^{2}_{e_{0}}(\omega)e^{i\omega t}.$ (40)
@@
and, following the steps of the derivation done previously for $\mathcal{C}$, the contributions ${\rm SFF}_{\text{d}}$ and ${\rm SFF}_{\mathrm{c}}$ are expressed as
@@
$\mathcal{C}_{\mathrm{c}}(t)\sim\int\text{d}t^{\prime}\mathcal{C}_{\text{d}}(t-t^{\prime}){\rm SFF}_{\mathrm{c}}(t^{\prime})\ .$ (44)
This can be interpreted as the ${\rm SFF}$ smoothed on the scale of the physical correlation function $\mathcal{C}_{\text{d}}$. Note that this is also valid at the level of the single instance, which therefore implies that the fluctuations intrinsic to the SFF also get smoothed out. At times $t\propto\mathcal{N}$, $\mathcal{C}_{\text{d}}(t-t^{\prime})$ acts as a delta function $\propto\delta(t-t^{\prime})$ for the slowly varying function ${\rm SFF}_{\mathrm{c}}$ and thus in the convolution one retrieves
$\mathcal{C}_{\mathrm{c}}(t)\sim f^{2}_{e_{0}}(0){\rm SFF}_{\mathrm{c}}(t).$ (45)
While the ${\rm SFF}$ and $\mathcal{C}$ seem to behave similarly at late times, it is not the case at early times when both functions are dominated by the slope. From Eqs. (36,42), it appears that $\mathcal{C}_{\text{d}}$ decays slower than ${\rm SFF}_{\text{d}}$ because of $f^{2}(\omega)$. This means that two different dip times are expected for $\mathcal{C}(t)$ and ${\rm SFF}(t)$.

### v4

@@ 1 Introduction
$\overline{\Gamma}(t)=\Gamma_{\text{d}}(t)+\Gamma_{\mathrm{c}}(t)\ ,$ (2)
for a suitably-defined connected correlation function $\Gamma(t)$ which depends on the RMT universality class of the system. We show that $\Gamma_{\text{d}}(t)$ amounts for the non-universal physical dynamical correlations (i.e. the Fourier transform of the smooth ETH off-diagonal function), while $\Gamma_{\mathrm{c}}(t)$ can be written as the convolution of the latter with the spectral correlations, and thus represents the spectral correlations responsible for the late-time ramp-plateau behaviour.
Similar observations have been drawn in the high-energy literature [7, 21, 22, 23, 24], random circuits [25] or hydrodynamics [26]. While the mechanism by which such a ramp can appear is easy to grasp, a justification of when it can be actually observed is instead rather subtle. Here, we provide a careful analytical study using the ETH approach. Our main assumption is that the average over energy levels and eigenvectors decouples. This is a standard result for rotationally-invariant Hamiltonians, and we here assume it also holds true for realistic many-body systems satisfying the ETH. We also show numerical evidence for our predictions in macroscopic many-body systems. While we chose to study numerically spin glass models, the same conclusions could have been drawn considering the SYK model, which however was not convenient for us as we need to treat separately different RMT universality classes. A key requirement to expect such a RMT behaviour is to consider observables whose expectation values (one-point functions) do not show any energy dependence. This allows, in absence of degenerate energy levels, the plateau to arise mainly from the fluctuations of the observable’s diagonal matrix elements, making it exponentially small in the system size. This is to be put in contrast with the case considered in [27, 28], and recently in [29], where an explicit energy dependence leads to a plateau which scales polynomially with the system size. Besides this in our calculations we assumed that the ETH function $f(\omega)$ which characterises the off-diagonal matrix elements has a finite value for $\omega\to 0$.
@@ 2.1 Two-time correlations and SFF in RMT
with the spectral form factor (SFF) defined as ${\rm SFF}(t)=\frac{1}{\mathcal{N}^{2}}\sum_{i,j}e^{i(E_{i}-E_{j})t}$. It is natural to introduce the shifted connected correlation function $\Gamma$ such that
$\Gamma(t)=C(t)-\left(1-\frac{\beta_{\rm RMT}}{2}\right)\lim_{t\to\infty}\overline{C}(t)\ ,$ (17)
@@
$\overline{\Gamma}(t)=\kappa_{2}\overline{{\rm SFF}}(t)\ .$ (18)
Note that the shift that we perform is necessary to observe the ramp, similarly to what happens in the partial spectral form factor [34]. This identity shows that, for a Hamiltonian drawn from random ensembles, the behaviour of a two-point function can be extracted from that of the ${\rm SFF}$ [7, 31]. The present analysis emphasises the importance of the symmetry classes for extracting the eigenvalue correlations from dynamical correlators, which is encoded in the re-scaled Eq.(17) and not in the bare connected correlator. From Eq.(18), it follows that, for random matrices, the decomposition of dynamical correlations in Eq.(2) holds upon identifying $\Gamma_{\text{d}}(t)=\kappa_{2}{\rm SFF}_{\text{d}}(t)$ and $\Gamma_{\mathrm{c}}(t)=\kappa_{2}{\rm SFF}_{\mathrm{c}}(t)$.
@@ 2.2 Lack of self-averaging
While the previous calculation characterised the ensemble-averaged quantity $\overline{\Gamma}(t)$, this section aims at estimating the single instance fluctuations $\delta\Gamma(t)=\Gamma(t)-\overline{\Gamma}(t)$. In random matrices, because $\overline{|A_{ij}|^{4}}\propto\overline{|A_{ij}|^{2}}^{2}$ with a proportionality constant depending on the ensemble, we expect that matrix elements behave as
@@
with $\{\xi_{ij}\}_{ij}$ a set of weakly correlated random variables with mean zero, variance one, and whose exact distributions depend on the observable considered. In a single instance $\Gamma(t)$, summing the $\mathcal{N}^{2}$ noise contributions gives rise to a noisy component of strength $\mathcal{N}^{-1}$, i.e.
$\Gamma(t)=\overline{\Gamma}(t)+\mathcal{O}(\mathcal{N}^{-1})\ .$ (20)
While this does not significantly obstruct the recovery of $\overline{\Gamma}(t)$ in the slope of order $\mathcal{O}(1)$ and the plateau of order $\mathcal{O}(\mathcal{N}^{-1})$, fluctuations completely hide the signal close to the dip which scales as $\mathcal{N}^{-3/2}$ (see Sec. 2.1).
This is confirmed by the numerics in the Gaussian ensemble case (see the right panels of Figs. 1-2 below) where we plot the single-sample $\Gamma(t)$. Contrary to the Spectral Form Factor, where the small-window time average reproduces the dip-ramp behaviour, for the dynamical correlation function $\Gamma(t)$, the dip and ramp are nowhere to be seen even after averaging over a small time window. The fact that, in the RMT toy models, the noise $\delta\Gamma(t)$ can become larger than the average signal $\overline{\Gamma}(t)$ thus results in correlation functions not being self-averaging, and in the small-window time-average not reproducing the ensemble-average in contrast to what is usually expected in ETH in the decaying part.
Figure 1: Left: $(a)$ Shifted correlation function $\overline{\Gamma}(t)$ and $(b)$ spectral form factor $\overline{{\rm SFF}}(t)$ averaged over 1000 GOE Hamiltonians. Right: $(c)$ Shifted correlation function $\Gamma(t)$ and $(d)$ spectral form factor ${\rm SFF}(t)$ obtained from a single GOE Hamiltonian. Plots $(c)$ and $(d)$ are for $L=12$ spins. In all plots, the thin curves are the exact result while the thick curves are a smoothed version obtained by convoluting with a Gaussian kernel. The horizontal dashed lines are the expected plateau values coming from the diagonal ensemble. The increasing oblique dashed lines are fits tproportional-$\propto t$.
@@ 2.3.1 The Gaussian Orthogonal Ensemble
We consider here a Hamiltonian from the GOE, i.e. drawn from Eq. (3) with $\beta_{\rm RMT}=1$, and take the observable $A=S^{z}_{j}$, with $j$ an arbitrary site index. Fig. 1 (right) shows the behaviour of $\Gamma(t)=C(t)-\kappa_{2}/\mathcal{N}$ and of the SFF for a single realisation of the GOE Hamiltonian. Upon averaging over some small time window, one cannot resolve the ramp in $\Gamma(t)$. However, if one considers the average $\overline{\Gamma}(t)$ over many realisations of the GOE Hamiltonian, the slope-dip-ramp-plateau becomes clear as shown in Fig. 1 (left). This is in contrast with the SFF where, despite the fluctuations, the ramp is visible also in a single instance, and the time average of one single instance coincides with the ensemble average.
@@ 2.3.2 The Gaussian Unitary ensemble
We now consider a Hamiltonian from the GUE (Eq. (3) with $\beta_{\rm RMT}=2$). Fig. 2 (right) shows the behaviour of $\Gamma=C(t)$ and of the SFF for a single realisation of the GUE Hamiltonian. Upon averaging over some small time window, the slope-dip-ramp-plateau feature becomes visible in the SFF, while the ramp cannot be resolved in $\Gamma$. If one averages $\Gamma$ and the SFF over many realisations of the GUE Hamiltonian, the ramp-plateau becomes much more clear as shown in Fig. 2 (left).
Figure 2: Left: $(a)$ Shifted correlation function $\overline{\Gamma}(t)$ and $(b)$ spectral form factor $\overline{{\rm SFF}}(t)$ averaged over 1000 GUE Hamiltonians. Right: $(c)$ Shifted correlation function $\Gamma(t)$ and $(d)$ spectral form factor ${\rm SFF}(t)$ obtained from a single GUE Hamiltonian. Plots $(c)$ and $(d)$ are for $L=12$ spins. In all plots, the thin curves are the exact result while the thick curves are a smoothed version obtained by convoluting with a Gaussian kernel. The horizontal dashed lines are the expected plateau values coming from the diagonal ensemble. The increasing oblique dashed lines are fits tproportional-$\propto t$.
@@ 2.3.3 Numerical evaluation of the fluctuations
In order to quantify the fluctuations of correlations relative to their average in Eq.(20), we consider the variance of $\Gamma(t)$ normalised by its average, namely
$\overline{\delta\Gamma^{2}}(t)/\overline{\Gamma}^{2}(t),$ (21)
@@
where $\delta{\rm SFF}(t)={\rm SFF}(t)-\overline{{\rm SFF}}(t)$ are the fluctuations of the SFF. The results for the GOE case are summarised in Fig. 3, and a similar picture can be drawn for the GUE case. While the ${\rm SFF}$ has a bounded noise-to-signal ratio, for the correlation function $\Gamma$, this ratio increases with the system size $L$ in the dip and ramp, signalling there that the fluctuations $\delta\Gamma(t)$ completely dwarf the ensemble average $\overline{\Gamma}(t)$.
@@
Figure 3: Normalised fluctuations of the shifted correlation function $\Gamma(t)$ and the spectral form factor ${\rm SFF}(t)$ averaged over 1000 GOE Hamiltonians at inverse temperature $\beta=0$. The thin curves are the exact result while the thick curves are a smoothed version obtained by convoluting with a Gaussian kernel.
@@ 3 Many-body Hamiltonian
$\overline{\Gamma}(t)=\Gamma_{\text{d}}(t)+\Gamma_{\mathrm{c}}(t)\ ,$ (23)
@@
$\Gamma_{\text{d}}(t)=\int\text{d}\omega\,|f_{e_{0}}(\omega)|^{2}e^{i\omega t}\ ,$ (24)
where $f_{e_{0}}(\omega)$ is the smooth function appearing in the off-diagonal ETH ansatz (see below). On the other hand, the second contribution $\Gamma_{\mathrm{c}}$ is related to the connected spectral correlations and generically features a late-time ramp followed by a plateau as for the SFF since
$\Gamma_{\mathrm{c}}(t)\sim\int\text{d}t^{\prime}C_{\text{d}}(t-t^{\prime}){\rm SFF}_{\mathrm{c}}(t^{\prime})\simeq f^{2}_{e_{0}}(0){\rm SFF}_{\mathrm{c}}(t)\ ,$ (25)
where we assumed that the function $f_{e}(\omega)$ has a finite limit for $\omega\to 0$.
@@ 3.1 Two-time correlations and SFF within ETH
An important assumption made in the following is that the observable $A$ has an average ${\mathcal{A}}(e)$ that does not depend on $e$, while $A$ still fluctuates as in Eq. (29). In fact, if $\partial_{e}\mathcal{A}(e)$, with $e=E/N$ the energy density, is of order one in the system’s size, there will always be polynomial corrections in $L$ coming from the diagonal part of the matrix elements after integration by saddle-point (see below). This hypothesis is required to avoid extra terms which could hide the slope-dip-ramp-plateau behaviour we are looking for [29], because of polynomially small/large corrections to the plateau, which may be understood by the fact that the fluctuations of the observable are then driven by the large local excursions of energy, see e.g. Ref.[29]. The assumption of no energy dependency is, for instance, guaranteed in the presence of symmetries and conservation laws. In particular one can study a local observable protected by some global conservation law in a disordered system, as will do in the following [^1] To be precise, we will focus on the following mechanism. We consider a disordered spin system that conserves the total magnetization $M^{z}=\sum_{i}\sigma_{i}^{z}$ and restricts ourselves to one of the sectors of the Hilbert space with fixed total magnetization $M\in\mathbb{Z}$. In the restricted Hilbert space, the thermal average of the total magnetization is, of course, $\langle M^{z}\rangle_{\beta}=M$. Then, we consider a given site $i$ and its local magnetisation $\langle\sigma_{i}^{z}\rangle_{\beta}$ which fluctuates depending on the disorder realisation. However, upon ensemble averaging, site-permutation symmetry is restored and $\overline{\langle\sigma^{z}_{i}\rangle_{\beta}}=M/L$. Therefore this quantity does not depend on temperature and energy.. Other examples could be drawn from disordered systems in their paramagnetic phase and at high energies where the ensemble average due to disorder implies the vanishing of expectation values of several quantities, at least in a certain range of temperatures (for instance the (mixed) $p$-spin model in transverse field), or from Floquet systems.
Drawing inspiration from the previous section on RMT, we now wish to establish a link between the large-time behaviour of connected correlation functions defined in Eq. (13) and the ${\rm SFF}$. To simplify the analysis, we consider the system at infinite temperature ($\beta=0$), since the extension to finite $\beta$ is a natural generalization. In fact, even at finite temperatures, energy-independent observables must still be considered to observe the exponentially small ramp, ensuring a simple generalization.
In the limit of large, yet finite, system size, the ensemble-averaged correlation function reads
@@
where $E=(E_{1}+E_{2})/2$ and $\omega=E_{1}-E_{2}$. Due to the standard saddle-point argument of the ETH, the contribution of the canonical average $\langle A\rangle^{2}$ exactly cancels that of the microcanonical average $\mathcal{A}(e)$ provided that $\partial_{e}\mathcal{A}(e)$ does not depend on energy on the saddle-point, as discussed above. Shuffling terms between both integrals gives
@@
$\Gamma(t)=C(t)-\left(1-\frac{\beta_{\rm RMT}}{2}\right)\lim_{t\to\infty}C(t)\ .$ (33)
@@
with $\frac{1}{L}\frac{\text{d}^{2}s}{\text{d}e^{2}}=-\beta^{2}/C_{V}$ and $C_{V}$ the heat capacity. This gives a contribution $\Gamma_{\text{d}}$ to the shifted correlation function $\Gamma$ defined by
$\Gamma_{\text{d}}(t)=\frac{1}{Z}\int\text{d}E\,\text{d}\omega\,\overline{\rho(E)}e^{\frac{1}{4L}\frac{\text{d}^{2}s(e)}{\text{d}e^{2}}\omega^{2}}f^{2}_{e}(\omega)e^{i\omega t}\ .$ (35)
As in the RMT case, $\Gamma_{\text{d}}$ is expected to give rise to the slope since it comes from the disconnected part of the correlations between energy levels. It is furthermore expected to be self-averaging with respect to the ensemble average, as is usual within the ETH. In the large system size limit, the factor $\exp\big{(}\frac{1}{4L}\frac{\text{d}^{2}s(e)}{\text{d}e^{2}}\omega^{2}\big{)}$ can be dropped and a saddle point approximation (also done for the partition function $Z$) leads to
${\Gamma}_{\text{d}}(t)=\int\text{d}\omega\,f^{2}_{e_{0}}(\omega)e^{i\omega t}\ ,$ (36)
where $e_{0}$ is such that $\frac{\text{d}s}{\text{d}e}(e_{0})=0$. Note that, if the observable were dependent on the energy, this quantity would be corrected by an additional term $\Delta_{e}^{2}(\partial_{e}{\mathcal{A}})^{2}$, with $\Delta_{e}^{2}\simeq N^{-1}$ the energy fluctuations, which would hide the ramp and the plateau, see e.g. Ref.[29].
@@
$\Gamma_{\mathrm{c}}(t)=$ $\frac{1}{Z}\int\text{d}E\,\text{d}\omega\,\pi^{2}\frac{\overline{\rho(E+\omega/2)}\,\overline{\rho(E-\omega/2)}}{\overline{\rho(E)}}\mathcal{R}\big{[}\mathcal{N}(\Phi(E+\omega/2)-\Phi(E-\omega/2))\big{]}f^{2}_{e}(\omega)e^{i\omega t}$
@@
$\Gamma_{\mathrm{c}}(t)\simeq\pi^{2}\int\text{d}\omega\,\mathcal{R}[\pi\overline{\rho(E_{0})}\omega]f^{2}_{e_{0}}(\omega)e^{i\omega t}.$ (40)
@@
and, following the steps of the derivation done previously for $\Gamma$, the contributions ${\rm SFF}_{\text{d}}$ and ${\rm SFF}_{\mathrm{c}}$ are expressed as
@@
$\Gamma_{\mathrm{c}}(t)\sim\int\text{d}t^{\prime}\Gamma_{\text{d}}(t-t^{\prime}){\rm SFF}_{\mathrm{c}}(t^{\prime})\ .$ (44)
This can be interpreted as the ${\rm SFF}$ smoothed on the scale of the physical correlation function $\Gamma_{\text{d}}$. Note that this is also valid at the level of the single instance, which therefore implies that the fluctuations intrinsic to the SFF also get smoothed out. At times $t\propto\mathcal{N}$, $\Gamma_{\text{d}}(t-t^{\prime})$ acts as a delta function $\propto\delta(t-t^{\prime})$ for the slowly varying function ${\rm SFF}_{\mathrm{c}}$ and thus in the convolution one retrieves
$\Gamma_{\mathrm{c}}(t)\sim f^{2}_{e_{0}}(0){\rm SFF}_{\mathrm{c}}(t).$ (45)
While the ${\rm SFF}$ and $\Gamma$ seem to behave similarly at late times, it is not the case at early times when both functions are dominated by the slope. From Eqs. (36,42), it appears that $\Gamma_{\text{d}}$ decays slower than ${\rm SFF}_{\text{d}}$ because of $f^{2}(\omega)$. This means that two different dip times are expected for $\Gamma(t)$ and ${\rm SFF}(t)$.